how to solve 25 th root of 33.86
Answer by Theo(13342)

i'm assuming you can use your calculator.

25th root of 33.86 is equal to 33.86 raised to the power of (1/25).

33.86^(1/25) = 1.151297279

to confirm that's correct, then raise 1.151297279 to the 25h power and you will get 33.86.

this assumes you used the stored answer in the calculator rather than entering the numbers again manually.

this is because the calculate stores about 15 decimal digits of accuracy but only display 8.

if you use the store number you get a more accurate answer.

using the stored answer, i got 1.121297279^25 = 33.86.

re-entering the displayed digits, i got 1.121297279^25 = 33.860000827.

it's close but no cigar because i was using a number rounded to 9 decimal digits rather than the internally stored number which contains more decimal digits than that.

the exponential form of a root is to raise the number to the reciprocal of the root.

in general, the nth root of x is equal to x^(1/n)

here's a reference on roots and exponents you might find helpful.

http://www.purplemath.com/modules/exponent5.htm